Financial accounting : record the impact of the following transations on the basic acounting equation and cashflows for a company that began operations on 1/1/18

1)Provided services and recived cash -3,000 6

2)Paid 450 in advance to rent a storage space for 6 months

3) preformed 2,500 of servies on account

4) accrued 800 of wages to be paid later

5)received 1,200 in advance for a one year service revenue contract

6)used cash to purchase equipment :cost - 2000; salavge value 2000 , useful life (6years)

7) adjust the records to record he expirtion of 4 months of rent on the storage unit (step#2)

8)earned 2 months of the service contract (step#5) record the adjusting entry

9)adjust the record to record 4 months of depreciation on the equipment

10)collected 1,700 from customers on account

balance sheet Income statment
assets liabilities equity
cash acct receviable prepaid rent equipment accu dep) = accrued wages unearned revenue retained earnings revenue expense = net inc. Cash flow
1)
2)
Answer #1
balance sheet Income statment Cash flow
assets liabilities equity
Event cash acct receviable prepaid rent equipment accu dep = accrued wages unearned revenue retained earnings revenue expense = net inc.
1 $ 3,000 $ 3,000 $ 3,000 $ 3,000 $ 3,000 Operating Activity
2 $   (450) $    450 $       -   $       -   $   (450) Operating Activity
3 $    2,500 $ 2,500 $ 2,500 $ 2,500
4 $        800 $   (800) $    800 $   (800)
5 $    1,200 $    1,200 $       -   $       -   $ 1,200 Operating Activity
6 $ (2,000) $    2,000 $       -   $       -   $(2,000) Investing Activity
7 $      (300) $   (300) $    300 $   (300)
8 $      (200) $    200 $    200 $    200
9 $      (100) $   (100) $    100 $   (100)
10 $    1,700 $ (1,700) $       -   $       -   $ 1,700 Operating Activity
Adjust the records to record he expirtion of 4 months of rent on the storage unit (450/6 = 75) (75*4) $        300
Earned 2 months of the service contract (step#5) record the adjusting entry (1200/12 = 100) (100*2) $        200
Dear student, Salvage value is equal to cost of purchase (Entry 6). It means no depreciation expense. So I can consider salvage value 200.
Adjust the record to record 4 months of depreciation on the equipment ((2000-200)/6) = 300 (300*4/12) $        100